### v1

# Context parroting: A simple but tough-to-beat baseline for foundation models in scientific machine learning

Abstract
Recently-developed time series foundation models for scientific machine learning exhibit emergent abilities to predict physical systems. These abilities include zero-shot forecasting, in which a model forecasts future states of a system given only a short trajectory as context. Here, we show that foundation models applied to physical systems can give accurate predictions, but that they fail to develop meaningful representations of the underlying physics. Instead, foundation models often forecast by context parroting, a simple zero-shot forecasting strategy that copies directly from the context. As a result, a naive direct context parroting model scores higher than state-of-the-art time-series foundation models on predicting a diverse range of dynamical systems, at a tiny fraction of the computational cost. We draw a parallel between context parroting and induction heads, which explains why large language models trained on text can be repurposed for time series forecasting. Our dynamical systems perspective also ties the scaling between forecast accuracy and context length to the fractal dimension of the attractor, providing insight into the previously observed in-context neural scaling laws. Context parroting thus serves as a simple but tough-to-beat baseline for future time-series foundation models and can help identify in-context learning strategies beyond parroting.
1 Introduction
A key test of generalization in scientific machine learning (SciML) is the ability to forecast future states of a new physical system. Prior SciML approaches primarily focused on developing specialized forecasting models trained extensively on specific dynamical systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. However, the generality of these models is limited by the amount of system-specific data available, motivating the recent development of time-series foundation models [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41], which are trained on vast amounts of observed and simulated time series, and which can subsequently perform zero-shot forecasts for any time series—including those generated by previously-unseen dynamical systems. Interestingly, it was recently found that, when available historical data is limited, time-series foundation models outperform classical deep learning models in forecasting chaotic dynamical systems [42].
What mechanisms do time-series foundation models use to make zero-shot forecasts, and why they are effective for dynamical systems? It was recently observed that one such foundation model, Chronos [35], often employs an extremely simple strategy when forecasting chaotic systems [42]. The strategy, context parroting, scans the context for repeating motifs and copies the part of the context following the best-matching motif as its prediction (Fig. 1). This can be viewed as a kind of “in-context nearest neighbor” algorithm, which has zero trainable parameters and thus requires no training. So how good is context parroting as a zero-shot forecasting strategy on its own?
Here, we compare context parroting with three state-of-the-art time series foundation models: Chronos [35], TimesFM [37], and TimeMoE [39] on the challenging task of forecasting chaotic systems. We find that context parroting outperforms all three foundation models in both zero-shot forecast accuracy and inference cost, especially for longer context windows. Our results suggest that current time series foundation models fail to extract physical insight from dynamical systems, and that their apparent performance is largely due to benchmark tasks that may be solved by simple repetition.
Our main contributions are:
  1. 1.
    Introduce context parroting as a simple but tough-to-beat baseline for zero-shot forecasting of dynamical systems
  2. 2.
    Show that context parroting outperforms time-series foundation models in predicting chaotic systems
  3. 3.
    Explain the in-context neural scaling law between forecast accuracy and context length, linking the scaling coefficient to the fractal dimension of the underlying attractor
2 Related work
2.1 Foundation models for science
Foundation models have recently been introduced for many scientific machine-learning tasks [43], including partial differential equations [44, 45, 46, 47, 48, 49, 50], neuroscience [51, 52, 53], and weather forecasting [54, 55]. However, most of these foundation models remain a black box and haven’t provided interpretable insights into the underlying process. Here, we analyze context parroting as a simple mechanism coming out of time-series foundation models, noting its strengths and weaknesses as a zero-shot forecasting strategy. This strategy, and the insights gained here, can potentially be applied to other scientific tasks.
2.2 In-context neural scaling law
Neural scaling laws describe the relation between the performance of a neural network and certain resources, such as model size, data size, or the amount of compute [56, 57, 58, 59]. Such scaling laws allow practitioners to predict the performance of yet-to-be-trained models based on the available resources and allocate them optimally [60].
When applying LLMs to forecast dynamical systems, Liu et al. [61] recently observed an in-context neural scaling law, in which the test loss decreases with the context window following a power law. Here, we show that this in-context neural scaling law can be reproduced when using context parroting to predict dynamical systems, and the scaling coefficient can be linked to an invariant property of the underlying dynamic process (the fractal dimension of the chaotic attractor). This finding shows that neural scaling laws are intrinsically linked to invariant properties of the process generating the data, and the theory can potentially be generalized to other models and tasks (e.g., can we estimate the “fractal dimension” of language from the neural scaling law of LLMs?).
2.3 In-context learning and induction heads
Induction heads are computational circuits that naturally emerge in simple transformers through training, and they have been hypothesized to underlie a lot of the in-context learning ability of foundation models [62, 63, 64]. In its simplest form, an induction head copies from repeating tokens in the context to make predictions. For example, when presented with a token stream $[A][B]\dots[A]$, an induction head will output $[B]$ as the next token. Prior works train transformers on minimal Markov chain grammars, and find that, during pretraining, models learn to identify increasingly higher-order $k$-grams, with different attention heads specializing in copying, lookup, and aggregation [65, 66]. These works imply that pretraining enables models to learn conditional distributions, allowing them to represent sequence distributions seen in the context [67, 68, 69, 70].
There is a clear parallel between context parroting and induction heads: both are essentially copy-and-paste operations, with context parroting involving the matching of not just one but multiple tokens. In fact, it is easy to imagine context parroting emerging naturally from combining multiple induction heads. This parallel can potentially explain the unreasonable effectiveness of applying language models trained on text to time-series tasks without fine-tuning or prompt engineering [27, 29, 30, 31, 61]. The induction heads formed from training on natural language happen to be also effective for predicting time series and can be easily repurposed to implement strategies such as context parroting.
3 Context parroting as an effective strategy for zero-shot forecasting
Overview of context parroting. In this section we motivate and introduce our baseline: context parroting. It was first identified as a strategy used by Chronos to predict chaotic systems, such as the partially observed Lorenz system shown in Fig. 1. It was also shown that one can predict the forecast accuracy of Chronos based on the context trajectory [42]: If there is a motif in the context that closely matches what happened immediately before the prediction starts, then Chronos would do well. This suggests that Chronos heavily utilizes context parroting to forecast chaotic systems. The parroting mechanism also explains Chronos’s sensitivity to initial conditions and sampling rate.
Figure 1: Example of a foundation model forecasting chaotic dynamics with context parroting. Here, the foundation model (Chronos) was asked to predict the $x$ variable of the Lorenz system based on a short context trajectory with $512$ data points. Blue is the prediction and gray is the ground truth. Chronos produced an accurate prediction by simply looking for a motif in the context similar to the motif immediately preceding the prediction (highlighted in yellow) and copying the evolution following the matching motif (highlighted by pink boxes). We distill this context parroting strategy into Algorithm 1 and compare it against time-series foundation models (including Chronos itself).
Context parroting is a simple forecasting strategy in which a model uses the last $D$ tokens of its context to query the remaining context. For whatever context sequence most closely matches the query, the following $H$ subsequent tokens are copied and used as the forecast. During the matching process, we exclude the last $D$ motifs to avoid parroting too close to where the prediction starts. Framed in terms of induction heads, the query lookup acts as a copy head, the nearest-neighbor match is a selector, and the exact repetion is the aggregation operation [66]. We describe this approach in Algorithm 1.
Algorithm 1 Context Parroting
Input: Context trajectory $x_{1:t}=\left\{x_{1},\ldots,x_{t}\right\}$, embedding dimension $D$ and forecast length $h$.
Output: Forecast trajectory $x_{t+1:t+h}=\left\{x_{t+1},\ldots,x_{t+h}\right\}$.
1:for all length-$D$ motif $s$: $x_{s-D+1:s}$ in the context $x_{1:t-D}$ do
2:     compute the distance $d_{s}$ between motif $s$ and the last motif $x_{t-D+1:t}$
3:Find the best-matching motif, $s_{opt}$, with the smallest distance
4:Set the first $t-s_{opt}$ predicted points to be $x_{t+1:2t-s_{opt}}=x_{s_{opt}+1:t}$ and repeat until the forecast length $h$ is reached
Mathematical Formulation. Context parroting corresponds to a continuous $1$-nearest-neighbor search over sequences of length $D$ in the context of length $L$. It thus corresponds to a limit of a Nadaraya–Watson model of the time series,
$\hat{p}(\mathbf{y}\mid\mathbf{q})=\frac{\sum_{j=D}^{L-H}K_{\sigma}\left(\mathbf{q},\,\mathbf{x}_{j-(D-1):j}\right)\;K_{\sigma}\left(\mathbf{y},\,\mathbf{x}_{j+1:j+H}\right)}{\sum_{j=D}^{L-H}K_{\sigma}\left(\mathbf{q},\,\mathbf{x}_{j-(D-1):j}\right)},$ (1)
where the query $\mathbf{q}$ represents the length-$D$ motif immediately preceding the end of the context window. $\mathbf{y}$ represents a length-$H$ forecast of subsequent values. The forecast sequence $\mathbf{y}$ has probability $\hat{p}$ conditioned on the query. The symmetric kernel $K_{\sigma}(\mathbf{u},\mathbf{v})=\sigma^{-d}K\bigl{(}(\mathbf{u}-\mathbf{v})/h\sigma\bigr{)}$ has bandwidth $\sigma$ in dimension $d=D\cdot\dim(x_{t})$. Assuming mean-squared error as a distance function in sequence space, we use a Gaussian kernel
$K_{\sigma}(\mathbf{u},\mathbf{v})=\dfrac{1}{(2\pi\sigma^{2})^{d/2}}\exp\left(-\dfrac{\|\mathbf{u}-\mathbf{v}\|^{2}}{2\sigma^{2}}\right)$
In practice, the second kernel on $\mathbf{y}$ in Eq. 5 becomes a delta function, and so we write the conditional mean predictor
$\hat{\mathbf{y}}(\mathbf{q})=\sum_{j=D}^{L-H}w(\mathbf{q},\mathbf{x}_{j-(D-1):j})\;\mathbf{x}_{j+1:j+H},$ $\quad w(\mathbf{q},\mathbf{z})\equiv\frac{K_{\sigma}\left(\mathbf{q},\,\mathbf{z}\right)}{\sum_{j=D}^{L-H}K_{\sigma}\left(\mathbf{q},\,\mathbf{x}_{j-(D-1):j}\right)}.$ (2)
Context parroting corresponds to the $1$-nearest-neighbor limit $\sigma\to 0$.
Context parroting preserves attractor properties at long context lengths. In Appendix C.3, we derive the following proposition,
$\lim_{L\to\infty}\mathbb{E}_{p}\!\left[F(\mathbf{y})|\mathbf{q}\right]=\mathbb{E}_{\mu}\!\left[F(\mathbf{x})\right]$
where $L$ is the context length for an Nadaraya–Watson estimator $p$, $F(\mathbf{y})$ is an estimate from a forecast sequence $\mathbf{y}$ of a property $F$ of an ergodic dynamical system, which has an invariant value $\mathbb{E}_{\mu}\!\left[F(\mathbf{x})\right]$ when calculated over the full attractor with underlying measure $\mu$. The query $\mathbf{q}$ is an arbitrary sequence of consecutive timepoints from the dynamical system. This proposition states that, when the context is sufficiently long, context parroting of an ergodic system preserves invariant values of the underlying dynamics. Context parroting thus represents an effective baseline for dynamical systems forecasting, because, in the limit of long context, it will preserve global properties like conditional distributions of values, Lyapunov exponents, or entropy production rates.
Relationship to classical nonlinear forecasting. We show in Appendix C.2 that, in various limits, context parroting reduces to two classical algorithms from nonlinear dynamics: the simplex projection technique and the S-map algorithm [71, 72]. Both approaches have their foundations in Takens’ embedding theorem, which states that time-delayed observables derived from a nonlinear dynamical system will recover key geometric properties of the underlying attractor [73]. However, unlike context parroting, which looks for the best matching motif, simplex projection tries to identify multiple matching motifs in the context and computes a weighted average as its forecast. This makes simplex projection more sensitive to the choice of the embedding dimension (i.e., the length of the motif $D$), limiting the method to small embedding dimensions in practice [74].
4 Methods
Datasets. The dysts dataset provides a standardized benchmark of 135 low-dimensional chaotic systems, each defined by a set of ordinary differential equations between dimensionality three and six [75]. Every system is annotated with its largest Lyapunov exponent $\lambda$, an invariant characteristic of the underlying dynamics that quantifies the rate at which small perturbations grow over time. In chaotic systems, even minor errors rapidly compound over a characteristic timescale known as the Lyapunov time, defined as $\tau\equiv\lambda^{-1}$. As a result, highly chaotic systems (with small $\tau$) are especially challenging to forecast.
Models. For time-series foundation models, we select Chronosbase (200M parameters), its variant Chronos-Boltbase (205M parameters), Time-MoElarge (200M parameters), and TimesFM-2.0 (500M parameters). All of these models are pretrained on massive amounts of real-world time series data (hundreds of billions of data points), which are often complemented by synthetic data to improve generalization. The three models encompass a wide array of design choices. Time-MoE and TimesFM employ decoder-only architectures, whereas Chronos adopts an encoder-decoder architecture. Chronos and Time-MoE use pointwise tokenization, whereas TimesFM chose patch tokenization. Time-MoE and TimesFM by default give point forecasts, whereas Chronos is designed to provide probabilistic forecasts with uncertainty quantification. Thus, for Chronos and Chronos-Bolt, we use the median prediction when evaluating forecast errors. An important parameter for these foundation models is the maximum context length $L_{\max}$. Chronos has the shortest allowed context window at $512$ data points, TimesFM-2.0 can handle $2048$ data points in the context, and Time-MoE has the largest capacity at $L_{\max}=4096$.
Pipelines. To evaluate different models’ ability to zero-shot forecast dynamical systems, we generate a chaotic trajectory of length $10^{5}$ for each of the $135$ chaotic systems in dysts, with a granularity of $30$ data points per Lyapunov time. For a given context length $L$, we randomly pick a length-$L$ segment from the chaotic trajectory and provide it to the model as the context. The model’s task is to predict the next $300$ data points (equivalent to $10$ Lyapunov times) solely based on the context. We ask the model to make a univariate forecast on each dimension independently, which is then evaluated separately for each dimension. To obtain reliable statistics, we aggregate the results over all $135$ chaotic systems, all dimensions, and $20$ random initial conditions for each system.
Metrics. In line with previous research [76, 77, 75, 78], we assess forecasting performance using two key metrics.
Symmetric Mean Absolute Percentage Error (sMAPE).
$\text{sMAPE}(\mathbf{x},\hat{\mathbf{x}})\equiv 2\dfrac{100}{T}\sum_{t=1}^{T}\dfrac{\left|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\right|}{\left|\mathbf{x}_{t}\right|+\left|\hat{\mathbf{x}}_{t}\right|},$
where the sequence $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}$ denotes the ground truth, and $\hat{\mathbf{x}}_{1},\hat{\mathbf{x}}_{2},\dots,\hat{\mathbf{x}}_{T}$ are the corresponding predictions made by the model.
Valid Prediction Time (VPT). This metric identifies the latest time step $t_{f}$ before which the sMAPE remains below a predefined threshold $\epsilon$, as described in Vlachas et al. [79]. Formally:
$\text{VPT}\equiv\text{argmax}_{t_{f}}\{t_{f}|\text{sMAPE}(\mathbf{x}_{t},\hat{\mathbf{x}}_{t})<\epsilon,\;\forall t<t_{f}\}.$
We use $\epsilon=30$, consistent with prior work [79, 78, 42].
In the appendix, we also show benchmark results using Mean Square Error (MSE) and Mean Absolute Error (MAE), two other metrics commonly used in the time series literature.
5 Results
5.1 Context parroting versus foundation models
Figure 2: Context parroting outperforms foundation models in zero-shot forecasting. Left: Forecast error of each model as a function of the forecast horizon. The context length is set to $512$ for all models. Right: Valid prediction time of each model as a function of the context length. Dashed lines indicate context lengths exceeding the maximum context window of the model. All results are averaged over the $135$ chaotic systems in the dysts database, with $20$ trajectories from random initial conditions for each system.
Here, we compare context parroting and foundation models in their ability to predict chaotic dynamics. Figure 2 shows each model’s forecasting error and valid prediction time. It is clear that context parroting is better than all foundation models in both metrics. In Fig. 9, we show that this remains true when benchmarked against MSE and MAE. Among foundation models, Chronos is the best performer in predicting chaotic systems, which is not surprising given that it utilizes parroting as a main forecasting strategy. Chronos’s tendency to context parrot arises from its distinct architecture as a language model that implicitly quantizes time series. As a result, Chronos is trained using cross-entropy loss, which incentivizes preservation of k-gram frequencies, ensuring repetition of sequences in the output and incentivizing generating diverse forecast samples consistent with the dynamical system’s underlying measure. In contrast, TimeMoE and TimesFM are trained using mean squared error loss, and so these models lose diversity and forecast the mean at long forecast horizons. Some representative forecasts from the foundation models are shown in Fig. 8.
Figure 3: Context parroting is computationally more efficient than foundation models. There is generally a gap of five or six orders of magnitude between context parroting and foundation models. For each foundation model, context lengths from $64$ to the maximum context window are considered. All inferences are performed with CPUs, and the forecast horizon is fixed to $300$ steps.
Moreover, as we demonstrate in Fig. 3, the inference cost of context parroting is negligible compared to all foundation models (not to mention the substantial GPU time needed to pre-train foundation models). For example, there is an over six orders of magnitude gap between Chronos and context parroting for all context lengths. Combined with the fact that the performance of parroting is not sensitive to the choice of the embedding dimension $D$ (Fig. 4), these results establish context parroting as a simple but hard-to-beat baseline for zero-shot forecasting of dynamical systems.
Figure 4: Effect of the embedding dimension $D$ on the forecast accuracy of context parroting. Overall, the valid prediction time stays consistent over a wide range of embedding dimension $D$. For short context windows, there is a slight advantage to small $D$. For long context windows, larger embedding dimensions are marginally better. This observation suggests potential improvements in the future that choose $D$ adaptively based on factors such as context length.
Figure 5 further explores the effect of context length on forecast accuracy. We find that longer context windows generally lead to better performance for both context parroting and Chronos. However, the longest context length Chronos can effectively utilize is $512$ data points; anything beyond that will only increase inference cost without improving accuracy. This limit is determined by Chronos’s maximum context window chosen at pre-training. To be able to utilize longer context, Chronos must be retrained from scratch with much more data and compute. Context parroting, on the other hand, is happy to utilize context data of any length. Interestingly, Chronos outperforms context parroting on short contexts, which points to additional zero-shot learning strategies beyond parroting employed by Chronos. This is perhaps not surprising given that at short context length, the times series becomes effectively nonstationary, which is the strength of time-series foundation models. Moreover, even when restricted to parroting, Chronos can in principle dynamically choose the optimal embedding dimension $D$ for each individual time series, giving it an advantage over parroting algorithms with a fixed $D$. It would be interesting to explicitly identify the mechanisms that enable Chronos to outperform parroting in this regime.
Parrot
Refer to caption Chronos
Figure 5: Parroting can better utilize longer context data while Chronos does better for shorter contexts. Each line represents a different context length. The performance of Chronos saturates once the context length exceeds $512$ data points, whereas the accuracy of context parroting keeps improving for longer context windows. Here we set the embedding dimension $D=10$ for the parroting algorithm.
5.2 In-context neural scaling law
Liu et al. [61] reported an in-context neural scaling law for LLMs applied to learning dynamical systems, in which the one-step forecast error decreases algebraically with context length. However, it is unclear where this scaling law came from or why LLMs trained on text can be effective for time series without fine tuning. Here, we show that context parroting naturally gives rise to the same in-context scaling law and provides geometric insights into its origin. Given the similarity between parroting and the induction heads implemented by LLMs [62], the geometric explanation we develop next for context parroting can conceivably be applied to LLMs and explain the observations in Liu et al. [61].
The left panel in Fig. 6 reproduces the power law relation between one-step forecast error (measured by sMAPE) and context length for the parroting method. Why do longer context lengths enable better predictions of chaotic systems? This is because more context data allows the algorithm to find better matching motifs, and a closely matched motif allows the parroted sequence to shadow the ground truth for longer. The overlap between the matching motifs can be measured by their Euclidean distance. For length-$D$ motifs, this is equivalent to embed the context trajectory in a $D$-dimensional delay-embedded space (mapping $x_{s}$ to $x_{s-D+1:s}$) and find the distance between the embedded last context point $x_{t-D+1:t}$ and its nearest neighbor. The right panel in Fig. 6 shows the improving overlap explicitly, with the distance between the matching motifs decreasing algebraically with context length.
Figure 6: Scaling laws with context length. Left: One-step forecast error vs context length. The scaling follows a power law for all embedding dimensions considered. Smaller $D$ is more accurate here because the use of the one-step forecast error. Larger $D$ can be more accurate for longer forecasting horizons. Right: Euclidean distance between the last context motif $x_{t-D+1:t}$ and its nearest neighbor in the $D$-dimensional delay-embedded space as a function of context length. Again, the scaling follows a power law for all $D$. The forecast accuracy is directly tied to the embedding distance: Smaller distances translate into better predictions. As the number of data points is increased, it becomes more and more likely that a context motif will land in the vicinity of the last context motif, and the correlation dimension of the chaotic attractor determines the rate of approach. In principle, infinite context length should give context parroting infinite accuracy for any stationary and deterministic systems.
Why then does the minimum embedding distance follow a power law with context length, $\ell\propto L^{-\alpha}$? We can link the scaling coefficient $\alpha$ to the fractal dimension of the chaotic attractor. The correlation dimension $d_{\mathrm{cor}}$ of an attractor is defined as
$d_{\mathrm{cor}}\equiv\lim_{\epsilon,\epsilon^{\prime}\rightarrow 0^{+}}\frac{\ln\left[\frac{C(\epsilon)}{C\left(\epsilon^{\prime}\right)}\right]}{\ln\left(\frac{\epsilon}{\epsilon^{\prime}}\right)},$
where $C(\epsilon)$ is the number of point pairs in the attractor that are within a given radius $\epsilon$ from each other. In other words, if we plot $C(\epsilon)$ against $\epsilon$ on a log-log plot, $d_{\mathrm{cor}}$ would be the slope of the plotted line. Due to the ergodic property of chaotic attractors, the context trajectory can be seen as a random sample of the attractor. Longer context trajectory contains more samples, and the expected distance between two context points in a delay-embedded space decreases with context length as $L^{-1/d_{\mathrm{cor}}}$. For example, for a two-dimensional attractor, the distance between two random points on the attractor will decrease as $1/\sqrt{L}$. Correlation dimension thus measures the speed at which the minimum embedding distance between points on an attractor can be reduced by including more samples, and higher dimensions require more points to reduce the distance to the same extent. Mathematically, we thus expect $\alpha=1/d_{\mathrm{cor}}$. However, numerically it is challenging to accurately estimate the scaling coefficient $\alpha$ due to noise in the data. Nonetheless, in Fig. 7, we observe strong correlation between $d_{\mathrm{cor}}$ and $1/\alpha$, supporting our theoretical argument above.
Figure 7: Strong correlation between the inverse scaling coefficient $1/\alpha$ and the correlation dimension $d_{\mathrm{cor}}$ of the chaotic attractor. We estimate $\alpha$ from the $\ell$ vs $L$ plot obtained with context parroting at $D=10$, where $L$ ranges from $2^{6}$ to $2^{16}$. This is done separately for each individual systems. Thus, each dot represents one of the $135$ chaotic systems included in our benchmark. The Spearman rank-order correlation coefficient is $0.44\pm 0.08$ (bootstrapped standard error), and a linear regression with standard error range is underset (shaded interval). The dashed line corresponds to an exact $1:1$ scaling.
6 Conclusion and future directions
We find that a simple forecast strategy outperforms leading foundation models on dynamical systems forecasting, a critical task in scientific machine learning. This surprising finding exposes a limitation of current time-series foundation models as general-purpose time-series forecasters and highlights the need to further scale them or to fine-tune them for specific domains. It also suggests that accurately measuring the performance of foundation models can be difficult for scientific machine learning tasks, because simple, uninformative strategies like parroting can effectively game both short- and long-term accuracy metrics.
Finding a simple but tough-to-beat baseline for a challenging task can encourage rethinking of the status quo, motivating better model architectures and baselines [80]. For example, identifying zero-shot learning strategies beyond parroting can spur the development of next-generation foundation models and contribute to the debate on whether (or to what extent) large language models are stochastic parrots [81, 82, 83, 84]. Context parroting formalizes an explicit baseline to compare against in the time-series domain and can help discover beyond-parroting strategies.
An important future direction is to generalize context parroting to deal with non-stationary time series while keeping the simplicity and efficiency of the method. A promising avenue involves combining parroting with a probabilistic model, such as a Gaussian Process, to account for nonstationary trends. Additionally, the diversity of long-term forecasts can be improved by allowing stochastic selection among multiple candidate forecasts. Once generalized, the non-stationary parroting method can replace Naive and Seasonal Naive to serve as a more informative baseline for the zero-shot forecasting of general time series (weather, traffic, finance, etc.).
Acknowledgments and Disclosure of Funding
YZ acknowledges support from the Omidyar Fellowship and NSF DMS 2436231. WG was supported by NSF DMS 2436233 and NSF CMMI 2440490. This project has been made possible in part by Grant No. DAF2023-329596 from the Chan Zuckerberg Initiative DAF, an advised fund of Silicon Valley Community Foundation.
Appendix A Sample predictions from foundation models
Figure 8: Example forecasts on a chaotic system from foundation models. This is the same task as presented in Fig. 1 (predicting the $x$ variable of the Lorenz system based on a short context trajectory with $512$ data points). Chronos does extremely well with a parroting strategy. The other models perform comparatively poorly and all exhibit a tendency to underestimate the oscillations (e.g., by quickly converging towards the mean). This is a general trend that we consistently observe across different chaotic systems and initial conditions.
Appendix B Benchmarking with other metrics
Figure 9: Context parroting outperforms foundation models in zero-shot forecasting. Same setup as in Fig. 2, but with the forecast error measured by MSE (left) and MAE (right).
Appendix C Context parroting
C.1 Discrete-token parroting
For fully-discrete tokens, an $L^{th}$ order Markov chain fit to the context has the form
$p(\mathbf{y}|\mathbf{q})=\dfrac{\#\{j:(x_{j-(L-1):j}=\mathbf{q})\land(x_{j+1:j+H}=\mathbf{y})}{\sum_{y^{\prime}}\#\{i:(x_{j-(L-1):j}=\mathbf{q})\land(x_{j+1:j+H}=\mathbf{y^{\prime}})}$ (3)
where the overall context has length $C$, and the Markov chain conditions the $H<C$ future tokens on the $L<C$ preceding tokens. The index $j$ runs over all contiguous sequences of length $L+H$ in the context, $j\in\{L-1,L,...,C-H-2,C-H-1\}$. The vector $\mathbf{q}\in\mathbb{R}^{L}$ represents the query, and the vector $\mathbf{y}\in\mathbb{R}^{H}$ represents the prediction in response to this query. Eq. 3 simply counts the number of token sequences of length $H+L$ that start with a given sequence of $L$ query tokens. A maximum-likelihood estimator derived from this model always samples the highest-likelihood sequence $\mathbf{y}$,
$\hat{\mathbf{y}}_{\text{MLE}}(\mathbf{q})=\text{argmax}_{\mathbf{y}}\log p(\mathbf{y}|\mathbf{q})$
However, this estimator may be unstable due to the appearance of queries $\mathbf{q}$ not seen in the context, motivating the use of token smoothing, in which Eq. 3 is replaced by the distribution
$p(\mathbf{y}|\mathbf{q})=\dfrac{\#\{j:(x_{j-(L-1):j}=\mathbf{q})\land(x_{j+1:j+H}=\mathbf{y})+\alpha}{\sum_{\mathbf{y}^{\prime}}\left(\#\{i:(x_{j-(L-1):j}=\mathbf{q})\land(x_{j+1:j+H}=\mathbf{y^{\prime}})+\alpha\right)}$ (4)
with increasing values of the parameter $\alpha$ causing predictions to converge to a uniform sample over possible predictions $\mathbf{y}$. The parameter value $\alpha=0$ reduces to no smoothing, while $\alpha=0.5$ corresponds to the Jeffreys prior and $\alpha=1$ corresponds to Laplace’s rule of succession.
C.2 Continuous-token parroting
A more general time series model treats tokens as continuous-valued. Some time series foundation models like Chronos use quantile binning to discretize time series values, allowing the direct use of discrete-token architectures [35]. However, many time series models assume effective continuity in token values, and we favor a continuous formulation in order to highlight connections to dynamical systems theory.
To model continuous-valued tokens directly, we replace the discrete count in §C.1 with a kernel-weighted estimate over all past subsequences. Let $\{\mathbf{x}_{t}\}_{t=1}^{C}$ be a univariate or multivariate time series. For context length $L$ and prediction horizon $H$, the Nadaraya–Watson estimate of the conditional density is
$\hat{p}(\mathbf{y}\mid\mathbf{q})=\frac{\sum_{j=L}^{C-H}K_{h}\left(\mathbf{q},\,\mathbf{x}_{j-(L-1):j}\right)\;K_{h}\left(\mathbf{y},\,\mathbf{x}_{j+1:j+H}\right)}{\sum_{j=L}^{C-H}K_{h}\left(\mathbf{q},\,\mathbf{x}_{j-(L-1):j}\right)},$ (5)
where $K_{h}(\mathbf{u},\mathbf{v})=h^{-d}K\bigl{(}(\mathbf{u}-\mathbf{v})/h\bigr{)}$ is a symmetric kernel with bandwidth $h$ in dimension $d=L\cdot\dim(x_{t})$. Assuming mean-squared error as a distance function in sequence space, we use a Gaussian kernel
$K_{h}(\mathbf{u},\mathbf{v})=\dfrac{1}{(2\pi h^{2})^{d/2}}\exp\left(-\dfrac{\|\mathbf{u}-\mathbf{v}\|^{2}}{2h^{2}}\right)$
In practice, we drop the second kernel on $\mathbf{y}$ in Eq. 5 and write the conditional mean predictor
$\hat{\mathbf{y}}(\mathbf{q})=\sum_{j=L}^{C-H}w(\mathbf{q},\mathbf{x}_{j-(L-1):j})\;\mathbf{x}_{j+1:j+H}$ (6)
where we have isolated a term corresponding to the weight of each sequence,
$w(\mathbf{q},\mathbf{z})\equiv\frac{K_{h}\left(\mathbf{q},\,\mathbf{z}\right)}{\sum_{j=L}^{C-H}K_{h}\left(\mathbf{q},\,\mathbf{x}_{j-(L-1):j}\right)}.$
Nearest-neighbor and global average limits. The bandwidth $h$ plays the role of a smoothing parameter (analogous to $\alpha$ in Eq. 4). As $h\to 0$ the scheme approximates a single-nearest neighbor parrot, while as $h\to\infty$ it converges to a global average over all sequences.
Connection to attention. If one takes
$K(\mathbf{u},\mathbf{v})=\exp\bigl{(}\mathbf{u}^{\top}\mathbf{v}/\tau\bigr{)},$
then Eq. 6 recovers a simplified form of softmax-attention, with the temperature hyperparameter $\tau$ controlling smoothness. In this view, the continuous parroting scheme is a kernel-regression analogue of discrete $k$–gram smoothing [85].
k-nearest-neighbor limit. We define a set $\text{Top}k$ corresponding to a subset of the possible values of the index $j\in\{L,L+1,...,C-H-1,C-H\}$. The $k$ elements of $\text{Top}k$ correspond to the indices $j$ that produce the $k$ largest values of $w(\mathbf{q},\mathbf{x}_{j-(L-1):j})$ across all values of $j$. We compute a simple average of these $k$ closest matches
$\hat{\mathbf{y}}(\mathbf{q})=\frac{1}{k}\sum_{j\in\text{Top}k}w(\mathbf{q},\mathbf{x}_{j-(L-1):j})\;\mathbf{x}_{j+1:j+H}$ (7)
yielding a $k$–nearest-neighbors parroting scheme. As $k$ increases, this estimator interpolates between exact parroting ($k=1$) and global average ($k\to C$).
Simplex projection.
Simplex projection, a classical forecasting method in nonlinear dynamics, corresponds to the special case $H=1$ (single step prediction), $k=L+1$ in Eq. 7. The condition $k=L+1$ represents the minimal number of affinely independent neighbors needed to triangulate a point in an $L$-dimensional space [71].
In simplex projection, the query $\mathbf{q}$ is interpreted as a time-delay embedding of the time series observable $\mathbf{x}$. Takens’ theorem argues that, under mild conditions, a finite number of time delay embeddings of an observable drawn from a deterministic ergodic system will be diffeomorphic (smoothly mappable) to the full-state dynamics [73]. Because simplex projection uses only neighbor identities, and not absolute distances, to weight context points, a delay embedding is sufficient to calculate the appropriate weights.
S-map forecasts.
Another common nonlinear forecasting technique retains all terms in the sum Eq. 6 , but instead performs a nonlinear weighting of the form
$K_{\theta}(\mathbf{u},\mathbf{v})=\exp(-\theta\|\mathbf{u}-\mathbf{v}\|/\bar{d})$
where the scale parameter $\bar{d}$ is determined by the distribution of distances among queries and points in the context. In practice, this parameter is often set to the mean pairwise distance among all sequences of length $L$ in the context. The optimal value of the hyperparameter $\theta$ increases as the underlying dynamics become more strongly nonlinear [72]. We note that, in the classical formulation of the S-map, a locally-linear model is fit based on all sequences of length $L+H$ seen in the context, while here we use the Nadaraya–Watson estimator in order to emphasize connections to modern kernel regression.
C.3 Invariants of motion
For ergodic dynamical systems in continuous time, there exists a natural measure $\mu(\mathbf{x})$ such that, for certain observables $F(\mathbf{x})$, the following condition almost surely holds,
[…]
where the second equality arises from the Birkhoff ergodic theorem [86].
We use the following convention for expectation values of sequences and single tokens; the expectation $\mathbb{E}_{\mu}\!\left[\mathbf{x}_{t:t+T}\right]$ refers to the expected value of the sequence $\mathbf{x}_{t:t+T}$ given pointwise measure $\mu$. We note that, for deterministic dynamical systems, once a given point is sampled on the attractor with measure $\mu(\mathbf{x}_{t})$, subsequent points have delta function conditional probability on the first point. Thus, we use the convention $\mu(\mathbf{x}_{t})=\mu(\mathbf{x}_{t:t+T})$ and we use the measure to refer to both the probability of a given timepoint, or a sequence of arbitrary length originating from that timepoint.
Proposition. Under appropriate kernel conditions,
$\lim_{C\to\infty}\mathbb{E}_{p}\!\left[F(\mathbf{y})|\mathbf{q}\right]=\mathbb{E}_{\mu}\!\left[F(\mathbf{x})\right]$
where $C$ is the context length for a Nadaraya–Watson estimator $p$, $F(\mathbf{y})$ is an estimate on a sequence $\mathbf{y}$ of an invariant property of an ergodic dynamical system with measure $\mu$, and $\mathbf{q}$ is an arbitrary sequence of consecutive timepoints from the dynamical system. This proposition states that, when the context is sufficiently long, a Nadaraya–Watson estimator of an ergodic system preserves the invariant values of the underlying dynamics.
Derivation. We start with the definition of the dynamical average,
[…]
Inserting Eq. 5 into this expression,
[…]
We multiply both the numerator and denominator by $1/C$ and take the limit $C\to\infty$, in order to convert the summations to expectations,
[…]
where $\mathbf{x}_{\leftarrow}$ represents the first $L$ points of random lookback window of length $L+H$ sampled from the underlying dynamical system, while $\mathbf{x}_{\rightarrow}$ denotes the next $H$ timepoints. In practice, this corresponds to a time series of $L+H$ points generated by simulating the dynamics starting at a point on the attractor randomly-sampled according to the measure $\mu$.
If we take the limit $h\to 0$, then the kernel $K_{h}$ becomes a delta function, yielding
$\lim_{h\to 0}\lim_{C\to\infty}\mathbb{E}_{\mu}\!\left[F(\mathbf{y})|\mathbf{q}\right]=\mathbb{E}_{\mu}\!\left[F(\mathbf{x}_{\rightarrow})|\mathbf{x}_{\leftarrow}=\mathbf{q}\right]$
If the measure $\mu$ is ergodic, then the conditional expectation of an invariant $F$ given any query $\mathbf{q}$ is simply its unconditional expectation,
$\lim_{h\to 0}\lim_{C\to\infty}\mathbb{E}_{\mu}\!\left[F(\mathbf{y})|\mathbf{q}\right]=\mathbb{E}_{\mu}\!\left[F(\mathbf{x})\right]$
C.4 Scaling laws
For a time series $\mathbf{x}_{1:T}$ with autocorrelation given by
$\left|\text{Corr}(\mathbf{x}_{t},\mathbf{x}_{t+\tau})\right|\leq Ce^{-\alpha\tau},\quad\alpha>0$
the expected forecast error scales as
$\mathbb{E}\!\left[\|\hat{\mathbf{y}}-\mathbf{y}\|^{2}\right]\sim e^{-\alpha L},\quad L\to\infty.$
Thus, under exponential decay of correlations (mixing), the amount of information about future states in a length-$L$ context window saturates exponentially quickly [87]. Thus, forecasts derived from increasingly large context windows converge exponentially quickly to optimal conditional forecasts under the invariant measure $\mu$.
Under standard smoothness conditions [88, 89], the forecast error also exhibits a standard bias-variance tradeoff of the form
$\mathbb{E}\!\left[\|\hat{\mathbf{y}}-\mathbf{y}\|^{2}\right]=\mathcal{O}(h^{4})+\mathcal{O}\left(\dfrac{1}{Ch^{L+H}}\right).$
The optimal width of the kernel thus scales as,
$h_{\text{opt}}\sim C^{-1/(4+L+H)}$