### v2

# Context parroting: A simple but tough-to-beat baseline for foundation models in scientific machine learning

Abstract
Recent time-series foundation models exhibit strong abilities to predict physical systems. These abilities include zero-shot forecasting, in which a model forecasts future states of a system given only a short trajectory as context, without knowledge of the underlying physics. Here, we show that foundation models often forecast through a simple parroting strategy, and when they are not parroting they exhibit some shared failure modes such as converging to the mean. As a result, a naive context parroting model that copies directly from the context scores higher than leading time-series foundation models on predicting a diverse range of dynamical systems, including low-dimensional chaos, turbulence, coupled oscillators, and electrocardiograms—and at a tiny fraction of the computational cost. We draw a parallel between context parroting and induction heads, which explains recent works showing that large language models can often be repurposed for time series forecasting. Our dynamical systems perspective also ties the scaling between forecast accuracy and context length to the fractal dimension of the underlying chaotic attractor, providing insight into previously observed in-context neural scaling laws. By revealing the performance gaps and failure modes of current time-series foundation models, context parroting can guide the design of future foundation models and help identify in-context learning strategies beyond parroting.
1 Introduction
A key test of generalization in scientific machine learning (SciML) is zero-shot forecasting: the ability to forecast future states of a new physical system based on a short context trajectory. Prior SciML approaches primarily focus on developing specialized forecasting models trained specifically on the system that needs to be predicted [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. However, the generality of these models is limited by the amount of system-specific data available, motivating the recent development of time-series foundation models [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41], which are trained on vast amounts of observed and simulated time series from diverse domains, and which can subsequently perform zero-shot forecasts for any time series—including those generated by previously-unseen dynamical systems. Interestingly, it was recently found that, when available historical data is limited, time-series foundation models outperform classical deep learning models in forecasting chaotic dynamical systems [42].
What mechanisms do time-series foundation models use to make zero-shot forecasts, and why they are effective for dynamical systems not seen during pre-training? It was recently observed that one such foundation model, Chronos [35], often employs an extremely simple strategy when forecasting chaotic systems [42]. The strategy, context parroting, scans the context for nearly repeating motifs and copies the part of the context following the best-matching motif as its prediction (Fig.˜1). This can be viewed as a kind of “in-context nearest neighbor” algorithm, which is easy to implement during in-context computation [43]. How good is context parroting as a zero-shot forecasting strategy? By comparing it with existing foundation models, what can we learn about current models’ strengths and limitations?
Here, we compare context parroting with three state-of-the-art time-series foundation models: Chronos and Chronos Bolt [35], TimesFM [37], and Time-MoE [39] on the challenging task of forecasting chaotic systems. We find that context parroting outperforms all three foundation models in both zero-shot forecast accuracy and inference cost, especially for longer context windows. Our results suggest that current time-series foundation models do not fully utilize the information in the context data, and thus still have significant room for improvement when it comes to SciML tasks.
Our main contributions are:
  1. 1.
    Introduce context parroting as a simple but effective baseline for zero-shot forecasting of dynamical systems, which can guide the design of more informative benchmarks that cannot be solved by simple repetitions and help identify forecasting strategies beyond parroting
  2. 2.
    Show that context parroting outperforms time-series foundation models in predicting chaotic systems and reveal common failure modes of the foundation models, which can guide the design of better models in the future
  3. 3.
    Explain the in-context neural scaling law between forecast accuracy and context length, linking the scaling coefficient to the fractal dimension of the underlying chaotic attractor
2 Related work
Foundation models for science. Foundation models have recently been introduced for many scientific machine-learning tasks [44], including partial differential equations [45, 46, 47, 48, 49, 50, 51], neuroscience [52, 53, 54], and weather forecasting [55, 56]. However, most of these foundation models remain a black box, and they have not yet provided interpretable strategies for forecasting diverse physical and dynamical processes. Here, we analyze context parroting as a simple mechanism used by time-series foundation models, noting its strengths and weaknesses as a zero-shot forecasting strategy. This strategy, and the insights gained here, can potentially be applied to other scientific tasks.
In-context neural scaling laws. Neural scaling laws describe the relationship between the performance of a neural network and certain resources, such as model size, data size, or the amount of compute [57, 58, 59, 60]. Such scaling laws allow practitioners to predict the performance of yet-to-be-trained models based on the available resources and allocate them strategically to optimize compute-adjusted accuracy [61]. When applying LLMs to forecast dynamical systems, Liu et al. [62] recently observed an in-context neural scaling law, in which the test loss decreases with the context length following a power law. Here, we show that this in-context neural scaling law can be reproduced when using context parroting to predict dynamical systems, and the scaling coefficient can be linked to an invariant property of the underlying dynamic process (the fractal dimension of the chaotic attractor). This finding shows that neural scaling laws are intrinsically linked to invariant properties of the process generating the data, and the theory can potentially be generalized to other models and tasks (e.g., can we estimate the “fractal dimension” of a language from the neural scaling laws of LLMs?).
In-context learning and induction heads. Induction heads are computational circuits that naturally emerge in simple transformers through training, and they have been hypothesized to underlie much of the in-context learning ability of foundation models [63, 64, 65, 66]. In its simplest form, an induction head copies repeating tokens in the context to make predictions. For example, when presented with a token stream $A][B]\dots[A]$, an induction head will output $[B]$ as the next token. Prior works train transformers on minimal Markov chain grammars, and find that, during pretraining, models learn to identify increasingly higher-order $k$-grams, with different attention heads specializing in copying, lookup, and aggregation [67, 68]. These works imply that pretraining enables models to learn conditional distributions, allowing them to represent sequence distributions seen in the context [69, 70, 71, 72].
There is a clear parallel between context parroting and induction heads: both are essentially copy-and-paste operations, with context parroting involving the matching of not just one but multiple contiguous tokens. In fact, it is easy to imagine context parroting emerging naturally from combining multiple induction heads. This parallel can potentially explain the unreasonable effectiveness of applying language models trained on text to time-series tasks without fine-tuning or prompt engineering [27, 29, 30, 31, 62]. The induction heads formed from training on natural language happen to be also effective for predicting time series and can be easily repurposed to implement strategies such as context parroting.
3 Context parroting as a zero-shot forecasting strategy
Overview of context parroting. In this section we motivate and introduce our baseline: context parroting. It was inspired by recent observations that Chronos often predicts chaotic systems by copying directly from the context [42]. An example of Chronos using parroting to forecast a partially-observed Lorenz system is shown in Fig.˜1.
Figure 1: Example of a foundation model forecasting chaotic dynamics with context parroting. Here, the foundation model (Chronos) was asked to predict the $x$ variable of the Lorenz system based on a short context trajectory with $512$ data points. Blue is the prediction and gray is the ground truth. Chronos produced an accurate prediction by simply looking for a motif in the context similar to the motif immediately preceding the prediction (highlighted in yellow) and copying the evolution following the matching motif (highlighted by pink boxes). We distill this context parroting strategy into Algorithm˜1 and compare it against time-series foundation models (including Chronos itself).
On a high level, context parroting uses the last $D$ tokens of the context to query the remaining context. For whatever context sequence that most closely matches the query, the subsequent tokens in the context are copied and used as the forecast. Because the length of the motif $D$ can be seen as the number of delayed states in a delay embedding from the lens of Takens’ embedding theorem [73], we also refer to $D$ as the embedding dimension and will use the terms embedding dimension and query length interchangeably. Interpreting $D$ as the embedding dimension is convenient because context parroting can be seen as a nearest neighbor algorithm in the $D$-dimensional delay-embedded space. During the matching process, we exclude the last $D$ motifs to avoid parroting too close to where the prediction starts. Framed in terms of induction heads, the query lookup acts as a copy head, the nearest-neighbor match is a selector, and the exact repetition is the aggregation operation [68]. We provide a pseudocode for context parroting in Algorithm˜1.
Algorithm 1 Context Parroting
Input: Context trajectory $x_{1:L}=\left\{x_{1},\ldots,x_{L}\right\}$, embedding dimension $D$ (i.e., the length of the motif to match), and forecast length $H$.
Output: Forecast trajectory $x_{L+1:L+H}=\left\{x_{L+1},\ldots,x_{L+H}\right\}$.
1:for all length-$D$ motif $s$: $x_{s-D+1:s}$ in the context $x_{1:L-D}$ do
2:  compute the Euclidean distance $d_{s}$ between motif $s$ and the last motif $x_{L-D+1:L}$
3:Find the best-matching motif, $s_{opt}$, with the smallest Euclidean distance
4:Set the first $L-s_{opt}$ predicted points to be $x_{L+1:2L-s_{opt}}=x_{s_{opt}+1:L}$ and repeat until the forecast length $H$ is reached
Mathematical Formulation. Context parroting corresponds to a continuous $1$-nearest-neighbor search over sequences of length $D$ in the context of length $L$. It thus corresponds to a limit of a Nadaraya–Watson model of the time series,
$\hat{p}(\mathbf{y}\mid\mathbf{q})=\frac{\sum_{j=D}^{L-H}K_{\sigma}\left(\mathbf{q},\,\mathbf{x}_{j-(D-1):j}\right)\;K_{\sigma}\left(\mathbf{y},\,\mathbf{x}_{j+1:j+H}\right)}{\sum_{j=D}^{L-H}K_{\sigma}\left(\mathbf{q},\,\mathbf{x}_{j-(D-1):j}\right)},$ (1)
where the query $\mathbf{q}$ represents the length-$D$ motif immediately preceding the end of the context window. $\mathbf{y}$ represents a length-$H$ forecast of subsequent values. The forecast sequence $\mathbf{y}$ has probability $\hat{p}$ conditioned on the query. The symmetric kernel $K_{\sigma}(\mathbf{u},\mathbf{v})=\sigma^{-d}K\bigl((\mathbf{u}-\mathbf{v})/h\sigma\bigr)$ has bandwidth $\sigma$ in dimension $d=D\cdot\dim(x_{t})$. Assuming mean-squared error as a distance function in sequence space, we use a Gaussian kernel
$K_{\sigma}(\mathbf{u},\mathbf{v})=\dfrac{1}{(2\pi\sigma^{2})^{d/2}}\exp\left(-\dfrac{\|\mathbf{u}-\mathbf{v}\|^{2}}{2\sigma^{2}}\right)$
We set the second kernel on $\mathbf{y}$ in Eq. 1 to a delta function, in order to output predictions that exactly match sequences from the context, rather than nearby sequences in a least-squares sense. We write the conditional mean predictor
$\hat{\mathbf{y}}(\mathbf{q})=\sum_{j=D}^{L-H}w(\mathbf{q},\mathbf{x}_{j-(D-1):j})\;\mathbf{x}_{j+1:j+H},$ $\quad w(\mathbf{q},\mathbf{z})\equiv\frac{K_{\sigma}\left(\mathbf{q},\,\mathbf{z}\right)}{\sum_{j=D}^{L-H}K_{\sigma}\left(\mathbf{q},\,\mathbf{x}_{j-(D-1):j}\right)}.$ (2)
Context parroting corresponds to the $1$-nearest-neighbor limit $\sigma\to 0$.
Context parroting preserves attractor properties at long context lengths. In Appendix F.3, we derive the following proposition,
$\lim_{L\to\infty}\mathbb{E}_{p}\!\left[F(\mathbf{y})|\mathbf{q}\right]=\mathbb{E}_{\mu}\!\left[F(\mathbf{x})\right]$
where $L$ is the context length for an Nadaraya–Watson estimator $p$, $F(\mathbf{y})$ is an estimate from a forecast sequence $\mathbf{y}$ of a property $F$ of an ergodic dynamical system, which has an invariant value $\mathbb{E}_{\mu}\!\left[F(\mathbf{x})\right]$ when calculated over the full attractor with underlying measure $\mu$. The query $\mathbf{q}$ is an arbitrary sequence of consecutive timepoints from the dynamical system. This proposition states that, when the context is sufficiently long, context parroting of an ergodic system preserves invariant values of the underlying dynamics. Context parroting thus represents an effective baseline for dynamical systems forecasting, because, in the limit of long context, it will preserve global properties like conditional distributions of values, Lyapunov exponents, or entropy production rates.
Relationship to classical nonlinear forecasting methods. We show in Appendix F.2 that, in various limits, context parroting is equivalent to two classical algorithms from nonlinear dynamics: the simplex projection technique and the S-map algorithm [74, 75]. Both approaches have their foundations in Takens’ embedding theorem, which states that time-delayed low-dimensional observables derived from a nonlinear dynamical system can recover key geometric properties of the underlying high-dimensional attractor [73]. However, unlike context parroting, which looks for the best matching motif, simplex projection tries to identify multiple matching motifs in the context and computes a weighted average as its forecast. This can potentially make simplex projection more sensitive to the choice of the embedding dimension $D$, limiting the method to small embedding dimensions in practice [76]. Other than simplex projection and S-map, there also exists other zero-shot forecasting strategies from nonlinear dynamics. For example, the Farmer-Sidorowich method [77] looks at multiple nearest neighbors in the context and builds a local linear model to make forecasts. An interesting future direction is to compare these methods from nonlinear dynamics with context parroting and time-series foundation models, which might inspire new zero-shot forecasting strategies.
4 Methods
Datasets. The dysts dataset provides a standardized benchmark of 135 low-dimensional chaotic systems, each defined by a set of ordinary differential equations between dimensionality three and six [78]. The chaotic systems are drawn from different published papers and span fields such as neuroscience, climate science, fluid dynamics, and astrophysics. Every system is annotated with its largest Lyapunov exponent $\lambda$, an invariant characteristic of the underlying dynamics that quantifies the rate at which small perturbations grow over time. In chaotic systems, even minor errors rapidly compound over a characteristic timescale known as the Lyapunov time, defined as $\tau\equiv\lambda^{-1}$. To normalize the difficulty of predicting different chaotic systems (so results from the $135$ systems can be meaningfully compared and combined), we generate trajectory data with a fixed sampling rate of $30$ points per Lyapunov time, and also measure the forecast performance in terms of Lyapunov times. To show the relevance of our findings to a broad class of SciML tasks, later we also go beyond low-dimensional chaotic systems and simulated data by benchmarking on real datasets from ECG measurements and electronic circuits.
Models. For time-series foundation models, we select Chronosbase (200M parameters), its variant Chronos-Boltbase (205M parameters), Time-MoElarge (200M parameters), and TimesFM-2.0 (500M parameters). All of these models are pretrained on massive amounts of real-world time series data (hundreds of billions of data points), which are often complemented by synthetic data to improve generalization. The three models encompass a wide array of design choices. Time-MoE and TimesFM employ decoder-only architectures, whereas Chronos adopts an encoder-decoder architecture. Chronos and Time-MoE use pointwise tokenization, whereas TimesFM chose patch tokenization. Time-MoE and TimesFM by default give point forecasts, whereas Chronos is designed to provide probabilistic forecasts with uncertainty quantification. Thus, for Chronos and Chronos-Bolt, we use the median prediction when evaluating forecast errors. An important parameter for these foundation models is the maximum context length $L_{\max}$. Chronos has the shortest allowed context window at $512$ data points, TimesFM-2.0 can handle $2048$ data points, and Time-MoE has the largest capacity at $L_{\max}=4096$.
Pipelines. To evaluate different models’ ability to zero-shot forecast dynamical systems, we generate a chaotic trajectory of length $10^{5}$ for each of the $135$ chaotic systems in dysts, with a granularity of $30$ data points per Lyapunov time. Each trajectory is normalized to have zero mean and unit standard deviation. For a given context length $L$, we randomly pick a length-$L$ segment from the chaotic trajectory and provide it to the model as the context. The model’s task is to predict the next $300$ data points (equivalent to $10$ Lyapunov times) solely based on the context. We ask the model to make a univariate forecast on each dimension independently, which is then evaluated separately for each dimension. To obtain reliable statistics, we aggregate the results over all $135$ chaotic systems, all dimensions, and $20$ random initial conditions for each system.
Metrics. In line with previous research [79, 80, 78, 81], we assess forecasting performance using a diverse set of complementary metrics.
Symmetric Mean Absolute Percentage Error (sMAPE).
$\text{sMAPE}(\mathbf{x},\hat{\mathbf{x}})\equiv 2\dfrac{100}{T}\sum_{t=1}^{T}\dfrac{\left|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\right|}{\left|\mathbf{x}_{t}\right|+\left|\hat{\mathbf{x}}_{t}\right|},$
where the sequence $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}$ denotes the ground truth, and $\hat{\mathbf{x}}_{1},\hat{\mathbf{x}}_{2},\dots,\hat{\mathbf{x}}_{T}$ are the corresponding predictions made by the model. To provide some context to help interpret the sMAPE value, we note that predicting the mean of white noise would give you an sMAPE around $200$.
Valid Prediction Time (VPT). This metric identifies the latest time step $t_{f}$ before which the sMAPE remains below a predefined threshold $\epsilon$, as described in Vlachas et al. [82]. Formally:
$\text{VPT}\equiv\text{argmax}_{t_{f}}\{t_{f}|\text{sMAPE}(\mathbf{x}_{t},\hat{\mathbf{x}}_{t})<\epsilon,\;\forall t<t_{f}\}.$
We use $\epsilon=30$, consistent with prior work [82, 81, 42]. For chaotic systems, consistently achieving a VPT over one Lyapunov time is considered impressive [81, 42].
Other than sMAPE and VPT, we also show benchmark results using Mean Square Error (MSE) and Mean Absolute Error (MAE), two other metrics commonly used in the time series literature.
For chaotic dynamical systems, point forecasts will inevitably fail due to the exponential rate of error accumulation. It is thus equally important for a forecasting model to preserve the long-term statistical and geometric properties of the chaotic attractors, such as Lyapunov exponents and the attractor dimension. Here, we compare the structure of true and predicted attractors by calculating the KL Divergence between their distributions.
Kullback–Leibler Divergence between Attractors ($D_{\mathrm{stsp}}$).
$D_{\mathrm{stsp}}\equiv D_{\mathrm{KL}}(P\|Q)=\sum_{x\in\mathcal{X}}P(x)\log\frac{P(x)}{Q(x)},$
where $P$ and $Q$ represent the true and the predicted attractor, respectively. When estimating $D_{\mathrm{stsp}}$, we follow the methodology in [83, 84]. Specifically, we place Gaussian kernels at each point in the true and predicted trajectories and estimate the KL divergence between these Gaussian mixtures using a sampling-based approximation [85].
In the appendix, we also measure attractor reconstruction accuracy using Correlation Dimension. The correlation dimension non-parametrically estimates the fractal dimension from a time series, by calculating the scaling of the number of attractor points that fall within a given radius of each point.
5 Results
5.1 Context parroting versus foundation models
Figure 2: Context parroting outperforms foundation models in zero-shot forecasting for both short-term point-wise accuracy and long-term attractor reconstruction. Left: Forecast error of each model as a function of the forecast horizon. The context length is set to $512$ for all models. Right: KL Divergence between the predicted attractors and the true attractors (smaller is better). Solid lines represent mean and dashed lines represent median. All results are obtained from $135$ chaotic systems in the dysts database, with 2020 trajectories from random initial conditions for each system.
Here, we compare context parroting and foundation models in their ability to predict chaotic dynamics. Figure˜2 shows each model’s forecasting error (measured by sMAPE) as well as their accuracy in attractor reconstruction (measured by KL Divergence). It is clear that context parroting is better than all foundation models tested here in both metrics. In Fig.˜7, we show that this remains true when benchmarked against MSE and MAE. The results for correlation dimension accuracy are shown in Fig.˜8, where parroting and Chronos significantly outperform the rest of the foundation models. We also explore the effects of observational noise and data granularity on the forecast performance in Tables˜4, 6, 5, 7, 8, 9, 11, 10, 13 and 12. Over the wide range of noise intensity and data granularity we tested, parroting is always the best or the second best according to all metrics (VPT, MAE, MSE, correlation dimension accuracy, and attractor KL Divergence).
Among foundation models, Chronos is the best performer in predicting chaotic systems, which is not surprising given that it utilizes parroting as a main forecasting strategy [42]. Chronos’s tendency to context parrot arises from its distinct architecture as a language model that implicitly quantizes time series. As a result, Chronos is trained using cross-entropy loss, which incentivizes preservation of k-gram frequencies and encourages the generation of diverse forecast samples consistent with the dynamical system’s underlying measure. In contrast, TimeMoE and TimesFM are trained using mean squared error loss. As a result, these models lose diversity and forecast the mean at long forecast horizons (i.e., they tend to underestimate the oscillations). Some representative forecasts from the foundation models are shown in Fig.˜6, which shows “regressing-to-the-mean” as a common failure mode for many foundation models when forecasting chaotic systems.
Figure 3: Context parroting is computationally more efficient than foundation models. There is generally a gap of five or six orders of magnitude between context parroting and foundation models. For each foundation model, context lengths from $64$ to the maximum context window are considered. All inferences are performed with CPUs, and the forecast horizon is fixed to $300$ steps.
Moreover, as we demonstrate in Fig.˜3, the inference cost of context parroting is negligible compared to all foundation models (not to mention the substantial GPU time needed to pre-train foundation models). For example, there is an over six orders of magnitude computational gap between Chronos and context parroting for all context lengths. Combined with the fact that the performance of parroting is not sensitive to the choice of the embedding dimension $D$ (Fig.˜10), these results establish context parroting as a simple but effective baseline for zero-shot forecasting of dynamical systems.
Parrot
Refer to caption Chronos
Figure 4: Parroting can better utilize longer context data while Chronos does better for shorter contexts. Each line represents the forecast error for a different context length. The performance of Chronos saturates once the context length exceeds its designed upper limit of $512$ data points, whereas the accuracy of context parroting keeps improving for longer context windows. Here we set the embedding dimension $D=10$ for the parroting algorithm.
Figure˜4 further explores the effect of context length on forecast accuracy. We find that longer context windows generally lead to better performance for both context parroting and Chronos. However, the longest context length Chronos can effectively utilize is $512$ data points; anything beyond that will only increase inference cost without improving accuracy. This limit is determined by Chronos’s maximum context window chosen at pre-training. To be able to utilize longer context, Chronos must be retrained from scratch with much more data and compute. Context parroting, on the other hand, is happy to utilize context data of any length. In Fig.˜9, we show VPT as a function of the context length for context parroting and all four foundation models. Again, parroting is the clear winner for medium to long context lengths, easily reaching an average VPT of over $5$ Lyapunov time when given long enough context.
Interestingly, Chronos outperforms context parroting on short contexts, which points to additional zero-shot learning strategies beyond parroting employed by Chronos. This is perhaps not surprising given that at short context length, the time series becomes effectively nonstationary, which is the strength of time-series foundation models. For example, Chronos is great at continuing the local trend in the context, which can be a more effective strategy than parroting when the length of the context is limited. Moreover, even when restricted to parroting, Chronos can in principle dynamically choose the optimal embedding dimension $D$ for each individual time series, giving it an advantage over parroting algorithms with a fixed $D$. It would be interesting to explicitly identify the mechanisms that enable Chronos to outperform parroting in the short-context regime.
5.2 In-context neural scaling law
Liu et al. [62] reported an in-context neural scaling law for LLMs applied to dynamical systems, in which the one-step forecast error decreases algebraically with context length. However, it is unclear where this scaling law came from or why LLMs trained on text can be effective for time series without fine tuning. Here, we show that context parroting naturally gives rise to the same in-context scaling law and provides geometric insights into its origin. Given the similarity between parroting and the induction heads implemented by LLMs [64], the geometric explanation we develop next for context parroting can conceivably be applied to LLMs and explain the observations in Liu et al. [62].
Figure 5: Scaling laws with context length. Left: One-step forecast error versus context length. The scaling follows a power law for all embedding dimensions $D$ considered. Smaller $D$ is more accurate here because the use of the one-step forecast error. Larger $D$ can be more accurate for longer forecasting horizons. Right: Euclidean distance between the last context motif $x_{L-D+1:L}$ and its closest match, as a function of context length. Again, the scaling follows a power law for all $D$. The forecast accuracy is directly tied to the motif distance: Smaller distances translate into better predictions. As the number of context data points is increased, it becomes more and more likely that a context motif will land in the vicinity of the last context motif, and the correlation dimension of the chaotic attractor determines the rate of approach. In principle, infinite context length should give context parroting infinite accuracy for any deterministic system with a well-defined attractor.
The left panel in Fig.˜5 shows the power law relation between one-step forecast error (measured by sMAPE) and context length for the parroting method. Why do longer context lengths enable better predictions of chaotic systems? This is because more context data allows the algorithm to find better matching motifs, and a closely matched motif allows the parroted sequence to shadow the ground truth for longer. The overlap between the matching motifs can be measured by their Euclidean distance. For length-$D$ motifs, this is equivalent to embed the context trajectory in a $D$-dimensional delay-embedded space (i.e., mapping $x_{s}$ to $x_{s-D+1:s}$) and find the distance between the embedded last context point $x_{L-D+1:L}$ and its nearest embedded neighbor. The right panel in Fig.˜5 shows the improving overlap explicitly, with the distance between the matching motifs decreasing algebraically with context length. It is easy to see that embedding distance should map linearly to the expected one-step forecast error, so the power law for the embedding distance implies the power law for one-step forecast error, as confirmed by Fig.˜5.
Why then does the minimum embedding distance follow a power law with context length, $\ell\propto L^{-\alpha}$? We can link the scaling coefficient $\alpha$ to the fractal dimension of the chaotic attractor. The correlation dimension $d_{\mathrm{cor}}$ of an attractor is defined as
$d_{\mathrm{cor}}\equiv\lim_{\epsilon,\epsilon^{\prime}\rightarrow 0^{+}}\frac{\ln\left[\frac{C(\epsilon)}{C\left(\epsilon^{\prime}\right)}\right]}{\ln\left(\frac{\epsilon}{\epsilon^{\prime}}\right)},$
where $C(\epsilon)$ is the number of point pairs in the attractor that are within a given radius $\epsilon$ from each other. In other words, if we plot $C(\epsilon)$ against $\epsilon$ on a log-log plot, $d_{\mathrm{cor}}$ would be the slope of the plotted line. Due to the ergodic property of chaotic attractors, the context trajectory can be seen as a random sample of the attractor. Longer context trajectory contains more samples, and the expected distance between two context points in a delay-embedded space decreases with context length as $L^{-1/d_{\mathrm{cor}}}$. For example, for a two-dimensional attractor, the distance between two random points on the attractor will decrease as $1/\sqrt{L}$. Correlation dimension thus measures the speed at which the minimum embedding distance between points on an attractor can be reduced by including more samples, and higher dimensionality requires more points to reduce the distance to the same extent. Mathematically, we thus expect $\alpha=1/d_{\mathrm{cor}}$. A similar scaling law has been derived for the Farmer-Sidorowich forecasting method from the nonlinear dynamics community [77]. Despite the theoretical correspondence, however, numerically it is challenging to accurately estimate the scaling coefficient $\alpha$ due to noise in the data. Nonetheless, in Fig.˜11, we observe strong correlation between $d_{\mathrm{cor}}$ and $1/\alpha$, supporting our theoretical argument above.
5.3 SciML tasks beyond low-dimensional chaotic systems
So far we focused on low-dimensional chaotic systems from the dysts dataset, which enabled systematic comparison between different forecasting models with standardized benchmarks. Here, we show that parroting also outperforms foundation models on a broader class of SciML tasks, including real-world datasets of current scientific interest. Our datasets cover fields such as fluid dynamics (von Karman vortex street with Reynolds number $\mathrm{Re}=900$, decomposed into PCA modes), neuroscience (ECG recordings from the QT Database in PhysioNet), engineering ($28$ coupled electronic circuits measured experimentally, from [86]), and physics ($23$ Kuramoto oscillators coupled through frustrated and nonreciprocal interactions, recently studied in [87]). These are all high-dimensional systems, two generated from simulations and two measured in the real world. For the metrics, we use MAE and MSE to measure pointwise forecast accuracy, and KL Divergence to measure the accuracy in attractor reconstruction. The results are summarized below. Parroting is either the best or the second best in all experiments. Other metrics, such as valid prediction time and correlation dimension accuracy, give similar rankings.
Task Parrot Chronos Chronos Bolt TimesFM TimeMoE
[…]
Table 1: Performance comparison (MAE @ 50 steps) of forecasting models across SciML tasks. Bold = best, italic = second best. Shading highlights best (dark) and second best (light).
Task Parrot Chronos Chronos Bolt TimesFM TimeMoE
[…]
Table 2: Performance comparison (MSE @ 50 steps) of forecasting models across SciML tasks. Bold = best, italic = second best. Shading highlights best (dark) and second best (light).
Task Parrot Chronos Chronos Bolt TimesFM TimeMoE
[…]
Table 3: Performance comparison (KL Divergence between predicted and true attractors) of forecasting models across SciML tasks. Bold = best, italic = second best. Shading highlights best (dark) and second best (light).
6 Conclusion and future directions
We find that a simple forecast strategy—context parroting—outperforms leading foundation models on dynamical systems forecasting, a critical task in scientific machine learning. This surprising finding exposes a limitation of current time-series foundation models as general-purpose time-series forecasters and highlights the need to further scale them or to fine-tune them for specific domains. It also suggests that accurately measuring the performance of foundation models can be difficult for scientific machine learning tasks, because strategies like parroting can effectively game both short- and long-term accuracy metrics.
Finding a simple but effective baseline for a challenging task can encourage rethinking of the status quo, motivating the development of better model architectures [88]. For example, context parroting formalizes an explicit baseline to compare against in the time-series domain and can help discover beyond-parroting strategies. Identifying zero-shot learning strategies beyond parroting can spur the development of next-generation foundation models and contribute to the debate on whether (or to what extent) large language models are stochastic parrots [89, 90, 91, 92].
An interesting future direction is to generalize context parroting to deal with non-stationary time series while keeping the simplicity and efficiency of the method. A promising avenue involves combining parroting with a probabilistic model, such as a Gaussian Process, to account for nonstationary trends. Additionally, the diversity of long-term forecasts can be improved by allowing stochastic selection among multiple candidate forecasts. Once generalized, the non-stationary parroting method can replace Naive and Seasonal Naive to serve as a more informative baseline for the zero-shot forecasting of general time series (weather, traffic, finance, etc.).
Finally, we want to emphasize that we are not proposing to replace time-series foundation models with context parroting. Instead, the value of parroting is as a simple baseline that can reveal the gaps in current foundation models and guide the design of future ones. When foundation models under perform relative to context parroting, it reveals that they haven’t learned to fully utilize the context data. For example, a common failure mode we observed across a range of leading foundation models (TimesFM, TimeMoE, Chronos Bolt) is that they tend to underestimate oscillations in the dynamics and the predictions often quickly converge to the mean (Fig.˜6). Being aware of context parroting as a baseline also guides the design of more informative benchmarks in the future. For example, for dynamical systems in dysts, one can generate test context data that are impossible to parrot by using only initial conditions that lead to trajectories without nearly repeating motifs.
Acknowledgments and Disclosure of Funding
YZ acknowledges support from the Omidyar Fellowship and NSF DMS 2436231. WG was supported by NSF DMS 2436233 and NSF CMMI 2440490. This project has been made possible in part by Grant No. DAF2023-329596 from the Chan Zuckerberg Initiative DAF, an advised fund of Silicon Valley Community Foundation.
Appendix A Sample predictions from foundation models
Figure 6: Example forecasts on a chaotic system from foundation models reveal common failure modes. This is the same task as presented in Fig.˜1 (predicting the $x$ variable of the Lorenz system based on a short context trajectory with $512$ data points). Chronos does extremely well with a parroting strategy. The other models perform comparatively poorly and all exhibit a tendency to underestimate the oscillations (e.g., by quickly converging towards the mean). This is a general trend that we consistently observe across different chaotic systems and initial conditions.
Appendix B Benchmarking with other metrics
Figure 7: Context parroting outperforms foundation models in zero-shot forecasting. Same setup as in Fig.˜2, but with the forecast error measured by MSE (left) and MAE (right).
Figure 8: Correlation dimension accuracy for parroting and foundation models. Each point represents the predicted correlation dimension of a chaotic attractor by a model (median of 2020 predictions from random initial conditions). The accuracy is measured by the Spearman correlation $r$ between the $135$ predicted correlation dimensions and the true correlation dimensions.
Figure 9: Valid prediction time as a function of the context length. Dashed lines indicate context lengths exceeding the maximum context window of the model. Results are averaged over $135$ chaotic systems in the dysts database, with 2020 trajectories from random initial conditions for each system.
Appendix C Effects of embedding dimension $D$
Fig.˜10 investigates how the choice of the embedding dimension $D$ affects the performance of context parroting. Overall, the valid prediction time stays consistent over a wide range of embedding dimension $D$. For short context windows, there is a slight advantage to small $D$. For long context windows, larger embedding dimensions are marginally better. This observation suggests potential improvements in the future that choose $D$ adaptively based on factors such as context length.
Figure 10: Effect of the embedding dimension $D$ on the forecast accuracy of context parroting. Each curve represents a different context length. Results are averaged over $135$ chaotic systems in the dysts database, with $100$ trajectories from random initial conditions for each system.
Appendix D Predicting scaling coefficient with the attractor dimension
Figure 11: Inverse scaling coefficient $1/\alpha$ versus correlation dimension $d_{\mathrm{cor}}$ of the chaotic attractor. We estimate $\alpha$ from the $\ell$ vs $L$ plot obtained with context parroting at $D=10$, where $L$ ranges from $2^{6}$ to $2^{16}$. This is done separately for each chaotic system. Thus, each dot represents one of the $135$ chaotic systems included in our benchmark. The Spearman rank-order correlation coefficient is $0.51\pm 0.08$ (bootstrapped standard error), and a linear regression with standard error range is underset (shaded interval). The dashed line corresponds to an exact $1:1$ scaling.
Appendix E Effects of noise and sampling rate
We add Gaussian noise to normalized chaotic trajectories and repeat the original experiments (a noise level of 0.1 translates to $10\%$ perturbation on each data point on average). The results are consistent across different orders of magnitude in noise, and parroting is consistently the best or the second best in all experiments.
Noise level Parrot Chronos Chronos Bolt TimesFM TimeMoE
[…]
Table 4: Valid prediction time across noise levels (higher is better). Bold = best, italic = second best. Shading highlights best (dark) and second best (light).
Noise level Parrot Chronos Chronos Bolt TimesFM TimeMoE
[…]
Table 5: MAE @ 1 Lyapunov Time across noise levels (lower is better). Bold = best, italic = second best. Shading highlights best (dark) and second best (light).
Noise level Parrot Chronos Chronos Bolt TimesFM TimeMoE
[…]
Table 6: MSE @ 1 Lyapunov Time across noise levels (lower is better). Bold = best, italic = second best. Shading highlights best (dark) and second best (light).
Noise level Parrot Chronos Chronos Bolt TimesFM TimeMoE
$10^{-3}$ 0.73 0.85 0.52 0.58 0.52
$10^{-2}$ 0.63 0.77 0.51 0.55 0.45
$10^{-1}$ 0.59 0.57 0.37 0.49 0.16
Table 7: Correlation dimension accuracy (Spearman correlation) across noise levels (higher is better). Bold = best, italic = second best. Shading highlights best (dark) and second best (light).
Noise level Parrot Chronos Chronos Bolt TimesFM TimeMoE
[…]
Table 8: KL Divergence between predicted and true attractors across noise levels (lower is better). Bold = best, italic = second best. Shading highlights best (dark) and second best (light).
In the main text, we set an intermediate granularity of 30 points per Lyapunov time. Below we compare it with results obtained for granularities of 10 points per Lyapunov time and 50 points per Lyapunov time. Granularity does not strongly affect the results or relative model ranking. Parroting is either the best or the second best in all experiments. This makes sense, as we would expect changing granularity to have a similar effect as rescaling of time (although with bigger or smaller gaps between data points). For example, if we use finer granularity by a factor of 2, then we would need to double the context length to get the same lookback window.
Granularity Parrot Chronos Chronos Bolt TimesFM TimeMoE
[…]
Table 9: Valid prediction time across different granularities (higher is better). Bold = best, italic = second best.
Granularity Parrot Chronos Chronos Bolt TimesFM TimeMoE
[…]
Table 10: MAE @ 1 Lyapunov Time across different granularities (lower is better). Bold = best, italic = second best.
Granularity Parrot Chronos Chronos Bolt TimesFM TimeMoE
[…]
Table 11: MSE @ 1 Lyapunov Time across different granularities (lower is better). Bold = best, italic = second best.
Granularity Parrot Chronos Chronos Bolt TimesFM TimeMoE
10 0.87 0.82 0.34 0.39 0.36
30 0.80 0.85 0.52 0.58 0.52
50 0.89 0.86 0.41 0.60 0.56
Table 12: Correlation dimension accuracy (Spearman correlation) across different granularities (higher is better). Bold = best, italic = second best.
Granularity Parrot Chronos Chronos Bolt TimesFM TimeMoE
10 0.087±0.137 0.127±0.173 0.573±0.307 0.444±0.326 0.467±0.368
30 0.122±0.194 0.173±0.209 0.346±0.297 0.345±0.298 0.354±0.290
50 0.137±0.207 0.230±0.256 0.406±0.305 0.361±0.323 0.370±0.338
Table 13: KL Divergence between predicted and true attractors across different granularities (lower is better). Bold = best, italic = second best.
Appendix F Context parroting
F.1 Discrete-token parroting
For fully-discrete tokens, a $D^{th}$ order Markov chain fit to the context has the form
$p(\mathbf{y}|\mathbf{q})=\dfrac{\#\{j:(x_{j-(D-1):j}=\mathbf{q})\land(x_{j+1:j+H}=\mathbf{y})}{\sum_{y^{\prime}}\#\{i:(x_{j-(D-1):j}=\mathbf{q})\land(x_{j+1:j+H}=\mathbf{y^{\prime}})}$ (3)
where the overall context has length $L$, and the Markov chain conditions the $H<L$ future tokens on the $D<L$ preceding tokens. The index $j$ runs over all contiguous sequences of length $D+H$ in the context, $j\in\{D-1,D,...,L-H-2,L-H-1\}$. The vector $\mathbf{q}\in\mathbb{R}^{D}$ represents the query, and the vector $\mathbf{y}\in\mathbb{R}^{H}$ represents the prediction in response to this query. Eq. 3 simply counts the number of token sequences of length $D+H$ that start with a given sequence of $D$ query tokens. A maximum-likelihood estimator derived from this model always samples the highest-likelihood sequence $\mathbf{y}$,
$\hat{\mathbf{y}}_{\text{MLE}}(\mathbf{q})=\text{argmax}_{\mathbf{y}}\log p(\mathbf{y}|\mathbf{q})$
However, this estimator may be unstable due to the appearance of queries $\mathbf{q}$ not seen in the context, motivating the use of token smoothing, in which Eq. 3 is replaced by the distribution
$p(\mathbf{y}|\mathbf{q})=\dfrac{\#\{j:(x_{j-(D-1):j}=\mathbf{q})\land(x_{j+1:j+H}=\mathbf{y})+\alpha}{\sum_{\mathbf{y}^{\prime}}\left(\#\{i:(x_{j-(D-1):j}=\mathbf{q})\land(x_{j+1:j+H}=\mathbf{y^{\prime}})+\alpha\right)}$ (4)
with increasing values of the parameter $\alpha$ causing predictions to converge to a uniform sample over possible predictions $\mathbf{y}$. The parameter value $\alpha=0$ reduces to no smoothing, while $\alpha=0.5$ corresponds to the Jeffreys prior and $\alpha=1$ corresponds to Laplace’s rule of succession.
F.2 Continuous-token parroting
A more general time series model treats tokens as continuous-valued. Some time series foundation models like Chronos use binning to discretize time series values, allowing the direct use of discrete-token architectures [35]. However, many time series models assume effective continuity in token values, and we favor a continuous formulation in order to highlight connections to dynamical systems theory.
To model continuous‐valued tokens directly, we replace the discrete count in §F.1 with a kernel‐weighted estimate over all past subsequences. Let $\{\mathbf{x}_{t}\}$ denote a univariate or multivariate time series. For context length $L$ and prediction horizon $H$, the Nadaraya–Watson estimate of the conditional density is
$\hat{p}(\mathbf{y}\mid\mathbf{q})=\frac{\sum_{j=D}^{L-H}K_{h}\left(\mathbf{q},\,\mathbf{x}_{j-(D-1):j}\right)\;K_{h}\left(\mathbf{y},\,\mathbf{x}_{j+1:j+H}\right)}{\sum_{j=D}^{L-H}K_{h}\left(\mathbf{q},\,\mathbf{x}_{j-(D-1):j}\right)},$ (5)
where $K_{h}(\mathbf{u},\mathbf{v})=h^{-d}K\bigl((\mathbf{u}-\mathbf{v})/h\bigr)$ is a symmetric kernel with bandwidth $h$ in dimension $d=D\cdot\dim(\mathbf{x}_{t})$ for the first kernel, and $d=H\cdot\dim(\mathbf{x}_{t})$ for the second kernel. Assuming mean-squared error as a distance function in sequence space, we use a Gaussian kernel
$K_{h}(\mathbf{u},\mathbf{v})=\dfrac{1}{(2\pi h^{2})^{d/2}}\exp\left(-\dfrac{\|\mathbf{u}-\mathbf{v}\|^{2}}{2h^{2}}\right)$
In practice, we drop the second kernel on $\mathbf{y}$ in Eq. 5 in order to output a prediction that exactly matches sequences from the context, rather than nearby sequences in a least-squares sense. We thus write the conditional mean predictor
$\hat{\mathbf{y}}(\mathbf{q})=\sum_{j=D}^{L-H}w(\mathbf{q},\mathbf{x}_{j-(D-1):j})\;\mathbf{x}_{j+1:j+H}$ (6)
where we have isolated a term corresponding to the weight of each sequence,
$w(\mathbf{q},\mathbf{z})\equiv\frac{K_{h}\left(\mathbf{q},\,\mathbf{z}\right)}{\sum_{j=D}^{L-H}K_{h}\left(\mathbf{q},\,\mathbf{x}_{j-(D-1):j}\right)}.$
Nearest‐neighbor and global average limits. The bandwidth $h$ plays the role of a smoothing parameter (analogous to $\alpha$ in Eq. 4). As $h\to 0$ the scheme approximates a single-nearest neighbor parrot, while as $h\to\infty$ it converges to a global average over all sequences.
Connection to attention. If one takes
$K(\mathbf{u},\mathbf{v})=\exp\bigl(\mathbf{u}^{\top}\mathbf{v}/\tau\bigr),$
then Eq. 6 recovers a simplified form of softmax‐attention, with the temperature hyperparameter $\tau$ controlling smoothness. In this view, the continuous parroting scheme is a kernel‐regression analogue of discrete kk–gram smoothing [93].
k-nearest‐neighbor limit. We define a set $\text{Top}k$ corresponding to a subset of the possible values of the index $j\in\{D,D+1,...,L-H-1,L-H\}$. The $k$ elements of $\text{Top}k$ correspond to the indices $j$ that produce the $k$ largest values of $w(\mathbf{q},\mathbf{x}_{j-(D-1):j})$ across all values of $j$. We compute a simple average of these $k$ closest matches
$\hat{\mathbf{y}}(\mathbf{q})=\frac{1}{k}\sum_{j\in\text{Top}k}w(\mathbf{q},\mathbf{x}_{j-(D-1):j})\;\mathbf{x}_{j+1:j+H}$ (7)
yielding a kk–nearest-neighbors parroting scheme. As $k$ increases, this estimator interpolates between exact parroting ($k=1$) and global average ($k\to L$).
Simplex projection.
Simplex projection, a classical forecasting method in nonlinear dynamics, corresponds to the special case $H=1$ (single step prediction), $k=D+1$ in Eq. 7. The condition $k=D+1$ represents the minimal number of affinely independent neighbors needed to triangulate a point in a $D$-dimensional space [74].
In simplex projection, the query $\mathbf{q}$ is interpreted as a time-delay embedding of the time series observable $\mathbf{x}$. Takens’ theorem argues that, under mild conditions, a finite number of time delay embeddings of an observable drawn from a deterministic ergodic system will be diffeomorphic (smoothly mappable) to the full-state dynamics [73]. Because simplex projection uses only neighbor identities, and not absolute distances, to weight context points, a delay embedding is sufficient to calculate the appropriate weights.
S-map forecasts.
Another common nonlinear forecasting technique retains all terms in the sum Eq. 6 , but instead performs a nonlinear weighting of the form
$K_{\theta}(\mathbf{u},\mathbf{v})=\exp(-\theta\|\mathbf{u}-\mathbf{v}\|/\bar{d})$
where the scale parameter $\bar{d}$ is determined by the distribution of distances among queries and points in the context. In practice, this parameter is often set to the mean pairwise distance among all sequences of length $D$ in the context. The optimal value of the hyperparameter $\theta$ increases as the underlying dynamics become more strongly nonlinear [75]. We note that, in the classical formulation of the S-map, a locally-linear model is fit based on all sequences of length $D+H$ seen in the context, while here we use the Nadaraya–Watson estimator in order to emphasize connections to modern kernel regression.
F.3 Invariants of motion
For ergodic dynamical systems in continuous time, there exists a natural measure $\mu(\mathbf{x})$ such that, for certain observables $F(\mathbf{x})$, the following condition almost surely holds,
$\mathbb{E}_{\mu}\!\left[F\right]\equiv\lim_{T\to\infty}\dfrac{1}{T}\int_{0}^{T}F(\mathbf{x})=\int F(\mathbf{x})d\mu(\mathbf{x})=\text{constant}$
where the second equality arises from the Birkhoff ergodic theorem [94].
We use the following convention for expectation values of sequences and single tokens; the expectation $\mathbb{E}_{\mu}\!\left[\mathbf{x}_{t:t+T}\right]$ refers to the expected value of the sequence $\mathbf{x}_{t:t+T}$ given pointwise measure $\mu$. We note that, for deterministic dynamical systems, once a given point is sampled on the attractor with measure $\mu(\mathbf{x}_{t})$, subsequent points have delta function conditional probability on the first point. Thus, we use the convention $\mu(\mathbf{x}_{t})=\mu(\mathbf{x}_{t:t+T})$ and we use the measure to refer to both the probability of a given timepoint, or a sequence of arbitrary length originating from that timepoint.
Proposition. Under appropriate kernel conditions,
$\lim_{L\to\infty}\mathbb{E}_{p}\!\left[F(\mathbf{y})|\mathbf{q}\right]=\mathbb{E}_{\mu}\!\left[F(\mathbf{x})\right]$
where $L$ is the context length for a Nadaraya–Watson estimator $p$, $F(\mathbf{y})$ is an estimate on a sequence $\mathbf{y}$ of an invariant property of an ergodic dynamical system with measure $\mu$, and $\mathbf{q}$ is an arbitrary sequence of consecutive timepoints from the dynamical system. This proposition states that, when the context is sufficiently long, a Nadaraya–Watson estimator of an ergodic system preserves the invariant values of the underlying dynamics.
Derivation. We start with the definition of the dynamical average,
$\mathbb{E}_{\mu}\!\left[F\right]=\int F(\mathbf{x})d\mu(\mathbf{x})$
Inserting Eq. 5 into this expression,
$\mathbb{E}_{\mu}\!\left[F(\mathbf{y})|\mathbf{q}\right]=\frac{\sum_{j=D}^{L-H}K_{h}\left(\mathbf{q},\,\mathbf{x}_{j-(D-1):j}\right)\;\int F(\mathbf{y})K_{h}\left(\mathbf{y},\,\mathbf{x}_{j+1:j+H}\right)d\mu(\mathbf{y})}{\sum_{j=D}^{L-H}K_{h}\left(\mathbf{q},\,\mathbf{x}_{j-(D-1):j}\right)},$
We multiply both the numerator and denominator by $1/L$ and take the limit $L\to\infty$, in order to convert the summations to expectations,
$\lim_{L\to\infty}\mathbb{E}_{\mu}\!\left[F(\mathbf{y})|\mathbf{q}\right]=\frac{\mathbb{E}_{\mu}\!\left[K_{h}(\mathbf{q},\mathbf{x}_{\leftarrow})\int F(\mathbf{y})K_{h}\left(\mathbf{y},\mathbf{x}_{\rightarrow}\right)d\mu(\mathbf{y})\right]}{\mathbb{E}_{\mu}\!\left[K_{h}(\mathbf{q},\mathbf{x}_{\leftarrow})\right]},$
where $\mathbf{x}_{\leftarrow}$ represents the first $D$ points of random lookback window of length $D+H$ sampled from the underlying dynamical system, while $\mathbf{x}_{\rightarrow}$ denotes the next $H$ timepoints. In practice, this corresponds to a time series of $D+H$ points generated by simulating the dynamics starting at a point on the attractor randomly-sampled according to the measure $\mu$.
If we take the limit $h\to 0$ (exact matching), then the kernel $K_{h}$ becomes a delta function, yielding
$\lim_{h\to 0}\lim_{L\to\infty}\mathbb{E}_{\mu}\!\left[F(\mathbf{y})|\mathbf{q}\right]=\mathbb{E}_{\mu}\!\left[F(\mathbf{x}_{\rightarrow})|\mathbf{x}_{\leftarrow}=\mathbf{q}\right]$
If the measure $\mu$ is ergodic, then the conditional expectation of an invariant $F$ given any query $\mathbf{q}$ is simply its unconditional expectation,
$\lim_{h\to 0}\lim_{L\to\infty}\mathbb{E}_{\mu}\!\left[F(\mathbf{y})|\mathbf{q}\right]=\mathbb{E}_{\mu}\!\left[F(\mathbf{x})\right]$
F.4 Scaling laws limiting prediction of stochastic systems
For a stochastic time series $\mathbf{x}_{1:T}$ with autocorrelation given by
$\left|\text{Corr}(\mathbf{x}_{t},\mathbf{x}_{t+\tau})\right|\leq Ce^{-\alpha\tau},\quad\alpha>0$
with $C$ representing a proportionality constant, the expected mean squared error of a forecast scales as
$\mathbb{E}\!\left[\|\hat{\mathbf{y}}-\mathbf{y}\|^{2}\right]\sim e^{-\alpha L},\quad L\to\infty.$
Thus, under exponential decay of correlations (mixing), the amount of information about future states in a length-$L$ context window saturates exponentially quickly [95]. Thus, forecasts derived from increasingly large context windows converge exponentially quickly to optimal conditional forecasts under the invariant measure $\mu$.
Under standard smoothness conditions [96, 97], the forecast error also exhibits a standard bias-variance tradeoff of the form
$\mathbb{E}\!\left[\|\hat{\mathbf{y}}-\mathbf{y}\|^{2}\right]=\mathcal{O}(h^{4})+\mathcal{O}\left(\dfrac{1}{Ch^{L+H}}\right).$
The optimal width of the kernel thus scales as,
$h_{\text{opt}}\sim C^{-1/(4+L+H)}$